OpenStudy (anonymous):

(e^(6x))/(e^(6x) + 1)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Integration problem

OpenStudy (anonymous):

let u=e^(6x)+1 then what is du/dx

OpenStudy (anonymous):

6e^(6x) so, dx = 1/(6e^6x) du

OpenStudy (anonymous):

where is the 1 coming from

OpenStudy (anonymous):

du = 6e(6x) dx or dx = (1/(6e^(6x)) du

OpenStudy (mathmale):

It'd be helpful if you could present integration problems either in Equation Editor or by drawing them in the Draw utility, below.\[\int\limits_{}^{}\frac{ e ^{6x} dx }{ 1+e ^{6x} }\]

OpenStudy (anonymous):

remember d/dx of e^x= itself

OpenStudy (mathmale):

Try this please: let u = 1 + e^(6x). Find du/dx. Unfortunately, your 6e(6x) is not correct. See magepkr's advice, above. Solve this for du.

OpenStudy (mathmale):

Examples:\[\frac{ d }{ dx }e^x=e^x\]

hartnn (hartnn):

i think pat meant, \(\Large dx = \dfrac{1}{6e^{6x}}du\) which is correct

hartnn (hartnn):

pat, you now have everything needed for substitution. go ahead and tell us what u get?

OpenStudy (mathmale):

Examples:\[\frac{ d }{ dx }e ^{3x}=e ^{3x}*\frac{ d }{ dx }3x=??\]

OpenStudy (mathmale):

To summarize: Let u be defined as follows:\[u=1+e ^{6x}\]

OpenStudy (mathmale):

Then\[\frac{ du }{ dx }=e ^{6x}*6\]so that \[du=6e ^{6x}dx,\]

OpenStudy (anonymous):

1/6 ᶘ 1 /(e^u +1) du = 1/6 (tan^-1 (e^u) =?

OpenStudy (mathmale):

or \[\frac{ du }{ 6 }=e ^{6x}dx\]

OpenStudy (mathmale):

Substitute du/6 for e^(6x) dx in the numerator. Substitute u for \[1+e ^{6x}\]

OpenStudy (mathmale):

in the denom. What do y ou get from doing this, Pat?

OpenStudy (mathmale):